OpenStudy (anonymous):

Find the period of the function. y = -3 cos x

OpenStudy (anonymous):

The period of cosine function is 2π. That doesn't change unless the variable is multiplied by a number. For example if you had y = -3 cos (2x) or y = -3 cos (½x). In those cases the period is 2π divided by the coefficient of x. - For y = -3 cos (2x) the period is 2π ÷ 2 = π. - For y = -3 cos (½x) the period is (2π ÷ ½) = 4π

jimthompson5910 (jim_thompson5910):

restating what peachpi said the general cosine function is y = A*cos(Bx - C) + D where |A| is the amplitude T = 2pi/B is the period C/B is the phase shift. Positive means right, negative means left y = D is the midline

OpenStudy (anonymous):

but there is only cos x so would the period still be 2pi then

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh okay thank both of your explanations helped i wrote them down

jimthompson5910 (jim_thompson5910):

In your case, B = 1 making the period T = 2pi/B = 2pi/1 = 2pi